4. On 1 January 2018, a borrower arranged a $2,980,000 three-year 3% bond payable, with interest paid annually each 31 December. There was an upfront fee of $83,461, which was deducted from the cash proceeds of the bond on 1 January 2018. The bond was issued at par.

Required:

a) What net amount is received on 1 January 2018?

b) Calculate the effective interest rate associated with the loan.

c) Calculate the interest expense reported by the borrower for each of the three years. (It may be easier to prepare an amortization table, although the table is not a required part of the answer.)[CW1]
Answer #1

a.

Net Amount Received = 2,980,000 - 83,461

Net Amount Received = $2,896,539

b.

Annual Payment = 0.03 * 2980000 = 89400

Time Period = 3 years

Using TVM calculation,

YTM = [FV = 2980000, PV = 2896539, PMT = 89400, T = 3]

YTM = 4.01%

c.

Interest Expense = $89,400 per year